OpenStudy (monae):

Determine the number of terms and identify the coefficient of each term of the expression. 1. -9x 2.3h-8a 3.2a-3s-z

OpenStudy (anonymous):

1. There is one term and the coefficient is -9 2. There are two terms and the coefficient of h is 3, the coefficient of a is-8 3. There are three terms and the coefficients are: 2 for a, -3 for s and -1 for z. Do you see what's going on?

OpenStudy (monae):

Wow is it really that easy? Lol I really suck at math can u help me with some others try and understand?

OpenStudy (anonymous):

Sure. Become a fan ;p

OpenStudy (monae):

How do you do that i have to sign up?

OpenStudy (anonymous):

There should be a blue link next to my name. It says, "Become a fan."

OpenStudy (anonymous):

If you don't see it, refresh your screen. brb

OpenStudy (monae):

add like terms if possible and write the result with the exponents in descending order. -4m^7-2m^7... I got -2m^7 is that right?

OpenStudy (anonymous):

Are you supposed to add -4m^7 and -2m^7?

OpenStudy (anonymous):

because if that's the case, you have\[-4m^7+-2m^7=-6m^7\]

OpenStudy (monae):

I have no clue I typed it just as it asked lol. I thought that since ooo wait your right because 4 and 2 arnt like terms. ok so -6^7 ok so the next one is -7y^8 - 4y^7=?

OpenStudy (monae):

So woiuld it be

OpenStudy (monae):

would*.... 3y^1 ???

OpenStudy (monae):

or since there are no like terms would it be -4y^7 - 7y^8 since thats descending orded

OpenStudy (anonymous):

Like terms are terms that have the same letter. So 7x and -4x would be like terms because they both share 'x'. 7x and 4y are not like terms because x and y are different. When they ask you to sum like terms, they're asking you to do something like the following: 3x -7y +2x -y = (3x+2x) + (-7y+-y) = 5x -8y

OpenStudy (anonymous):

Can you type out the question?

OpenStudy (monae):

Add like terms, if possible, and write the result with the exponents in descending order. 1... -7y^8 - 4y^7 = .. so I figured it would be -4y^7 - -7y^8 seeing how I cant add the terms

OpenStudy (anonymous):

Nearly. Though, I'm glad you saw that you couldn't add the terms. The question asked for exponents in descending order, so you'd have to rearrange your answer as -7y^8-4y^7 which looks, to me, like what they gave you in the question anyway, by coincidence.

OpenStudy (monae):

o i had it backwards. ok. so here is another problem with the same question.... -3m^6+8m^5-4m^4+5m^6-2m^5 .... = 2m^6+6m^5-4m^4 ??

OpenStudy (anonymous):

lol, what happened to your m^5 term? What's 8 + 5?

OpenStudy (monae):

13... but since its -2m^5 wouldnt it be 8m^5-2m^5 to = 6m^5?

OpenStudy (anonymous):

Oh I'm sorry, I stuffed it. It's hard deciphering. You're absolutely right. You have the hang of it :)

OpenStudy (monae):

lol ok.. and i know it would make it easier if the powers actually worked on here so the numbers would be above,. So it says... Perform the indicated operation. 1. (-2 + 3n^7 - 8n^5) + (9n^7 + 6n^5 - 5).. i havent worked on this one yet so let me try it then tell me if i did anything wrong

OpenStudy (anonymous):

Okay. I'm glad you're trying them out before asking.

OpenStudy (monae):

Yea I have to get this down. I am not the type of person to just get answers because then it takes me no where in life

OpenStudy (monae):

The person typin is my lil sis rejanae

OpenStudy (anonymous):

\[(-2+3n^7-8n^5) + (9n^7 + 6n^5 - 5)\]

OpenStudy (anonymous):

lol told you it works

OpenStudy (anonymous):

You have an excellent attitude Monae.

OpenStudy (anonymous):

You'll go far.

OpenStudy (monae):

\[3+12n^7-2n^5 or is it 3+12n^14-2n^10 /]

OpenStudy (monae):

its suppose to be ^14 and ^10

OpenStudy (monae):

\[3+12n^7-2n^5\]

OpenStudy (monae):

or \[3+12n^14-2n^10\]

OpenStudy (anonymous):

okay, hang on, just need to check something.

OpenStudy (monae):

Lol ok. and Thank You

OpenStudy (anonymous):

It's\[12n^7-2n^5-7\]

OpenStudy (anonymous):

The 7 is actually\[7n^0\]I don't know if you've ever seen something like that. Whenever you raise a number to the power of 0, the result is 1. So when we order from highest to lowest powers, the constant goes on the end because the power of n is the lowest...0.

OpenStudy (monae):

These don't have to be in descending order for the exponents so the only thing i got wrong woul,d be the 3.

OpenStudy (anonymous):

Yeah, I really think you've got the hang of it. From here on, it's just practice.

OpenStudy (monae):

Yea the \[7^0\] is in chapter 5.6 lol i hvnt gotten there yet. im stil stuck on this lol

OpenStudy (monae):

So do I still have to write it descending like you did or is my answer right and I just have to replace the 3 with a 7

OpenStudy (anonymous):

If they don't ask for it, TECHNICALLY it's correct. BUT mathematicians, as a matter of course, always write the polynomial in descending order BECAUSE it makes our lives easier in the end when we're trying to calculate things. I suspect it would be easier for your teachers too.

OpenStudy (monae):

Only problem is that my proffesor doesn't check to see if you have the right answers he only checks to see if you have something written down and that you tried the problem. I think its very idiotic to have that seeing how we won't learn if we did the homework right or not.

OpenStudy (anonymous):

Yeah, I agree.

OpenStudy (anonymous):

That's not a very good way of teaching.

OpenStudy (anonymous):

But you're attitude is what will save you. You sorted this out in several minutes.

OpenStudy (monae):

Thank You soo much. I understand it now.. But trust I will be back on later for your help with those 7^0 and other problems like that.. will you be on tomorrow around the same time or friday anytime after 3?

OpenStudy (anonymous):

Maybe...you can always try. You might get an e-mail from my response here. If you click on 'view answer' or the link, it will take you back to this post. You can ask more questions here as a means of getting my attention, or you can ask others if you can't wait, by using the 'As a question...' box.